OpenStudy (anonymous):

A right triangle has a hypotenuse of length 25 and a leg of length 15. What is the length of the other leg? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

can some one please show me how to set it up

OpenStudy (anonymous):

Use the Pythagorean Theorem. a² + b² = c² 20² + b² = 25² 400 + b² = 625 b² = 625 - 400 b² = 225 (Square root) b = 15 (Answer) The other leg is 15. Hope this helps.